Tech Staff Meeting Notes May 3, 2016

The top priority is always repair of down stations.
NM problem stations
ET problem stations

Duty Table updated June 2014


News & Highlights

The Fish Fry will be on Friday May 6.

NetOpsVIII will be in Pasadena October 11-13. It will focus on field ops and be about 2.5 days. We should be able to send two or three people.

James has successfully rebooted the SIS effort. We discussed the pros and cons of defining the nodes as multistation digitizers, and just faking it and lumping the entire response into a single piece of hardware for the analog stations. The consensus was we don't need to track analog hardware in SIS so we'll fake it.


2016 Preventive Maintenance 0% complete
The 2015 PM's were at about 87% at the end of the year. Most of the remaining stations were SM stations in relatively civil environments. Need to check and clear the flash on them.


Taken from:


Misc Problems
CPAR - spikes
HOVM - ditto
TYAR - ditto
SEAR.HNE -need quake to verify

MORT was flooded. Remains inaccessible.

All ISIS3 hardware except that slated for PLAR is installed. Have about 26 remaining, mostly MKTA and some NMAD outer stations.

There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.

NODE Tasks

All nodes are due a firmware upgrade that will sufficiently change things, the the initial upgrade and configure should be done at CERI, then swap hardware at the node.

John checked other networks one plans to replace Windows digiziting boxes and found none. NC has migrated to the NI X-series boards and adsend does work with it though its still on MS Windows. John will try to find out which specific and OS is used.

DBB: Move the DBB link at CERI from the west end of 3892 to the east side of 3904 and eliminate the thin-wire.


LNXT: Generator failure on May 1. There was a power outage over the weekend and the generator burned an excessive amount of oil causing it to fail. Additionally, the transfer switch did not switch back to Line when it came back. Dave will write a document describing what happened, as we know it, what failed and needs to be tested and repaired, and recommended remedial actions.

Uverse is likely avaialable at the lnxt node. Steve will investigate to see if its cost effective to upgrade from dsl. We don't need the band width but service for dsl will continue to deteriorate. Current costs for lnxt dsl is $119.55/month.




Backup Liebert has a bad blower. Workorder submitted

Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave.


Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to add isolation mounts and test farther away. We had to return the compact so will need to use a 6TD instead. We need this cell modem back by June. Noise test to begin tomorrow 5/4.


Taken from:


Misc Problems

GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.

John's updated todo list .


Improve RCGA AC power and perform maintenance work on shed.

Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are t oo tall to get a signal out. Looking for an alternate site.


HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established. Still waiting on completion of communcations to get on local internet. We need this modem back before June 30. John has a Thursday 5/5 meeting to do this.

Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.

GFM - Need to do a noise study for a closer site, then move it in the spring along with upgrading to S-13.


Replace VHTN E-W and N-S Instruments.

WMTN pit needs to be moved due to erosion.

We're still repeating CPRT through GTTN. The new internet host at the Corryton Senior Center for CPRT has been formally approved by Greg Hill the center coordinator and the Knox County facility supervisor.


Bob has funding for a new CUS station. Deadline is Aug 31 for obligation of all funds. It will be a drill braced monument rather than the piling style. Siting should be in the Hayti or Blythville area. If Blythville, it needs to be to the south of the seismicity (e.g. south of the Cottonwood Grove Fault, the line of seismicity that runs between Marked Tree and south of Dyersberg). Chris will look for possible sites and Bob will work on 3 bids for the monument installation.

CVMS has intermittent connection problems.

HCES needs an antenna swap.

AG Network

TA/AO adoptions. Both stations have had initial visits, are dry, and look to be a go. Need to get refteks and cabling from ASL to replace the Quantaras. Then make the switch sometime in June. At a future date, we'll reinstall as much hardware as possible into a surface enclosure. These will be AG stations if AGS is able to obtain O&M funds.

FCAR AC power upgrade was abandoned. A tree was removed improving sunlight Temporay additional solar was installed and seems to be working. Need to make it permanent. Done.

HHAR got its annual PM.


Needs some clean up.

The FDL wiki is here:


The seisnet group on umdrive is at You will need your uuid and password.

The seisnet wiki is at You will need your uuid and password.